solve system of equations using multiplication with the addition method;
7x-3y=-5
3x+2y=11

Answer by mananth(16949)

7x-3y=-5....................1
3x+2y=11....................2
..
multiply (1) by 2 & (2) by 3
14x-6y=-10
9x+6y=33
..
add them up
14x+9x=-10+33
23x=23
/23
x=1
..
plug value of x in (1)
7-3y=-5
3y=7+5
3y=12
/3
y=4
...
m.ananth@hotmail.ca